# Inner products of deep level Deligne–Lusztig representations of Coxeter type

### v1

@@ 3.2. An extension of action
We define an action of $\mathbb{T}^{F}\times\mathbb{T}^{F}$ on $\Sigma_{w}^{i}$ by
@@
$H_{w,\chi}=\{(t,t^{\prime})\in\mathbb{T}_{h}\times\mathbb{T}_{h};w^{-1}t^{-1}F(t)w={t^{\prime}}^{-1}F(t^{\prime})\in\operatorname{Im}(\chi)\}.$
@@
$N_{F}^{F^{N_{0}}}:\mathbb{T}_{h}\mathbb{T}_{h},\quad ttF(t)\cdots F^{N_{0}-1}(t).$
@@ 3.3. Handling $\Sigma_{w}^{0}$ for $w\in W_{e}\setminus W_{e}^{c\sigma}$
$(\hat{\Sigma}_{w,u}^{i})^{\mathbb{G}_{m}}\subseteq\{1\}\times\{1\}\times\{1\}\times\mathbb{T}\times\{1\}\times\{1\}.$
@@ 3.4. Handling $\Sigma_{w}^{0}$ for $w\in W_{e}^{c\sigma}$
${}^{\flat}\Sigma_{w}^{i}=\{(x,x^{\prime},y)\in(F\mathbb{U}\cap\mathbb{U}^{-})\times F^{i}(F\mathbb{U}\cap\mathbb{U}^{-})\times\mathbb{G};xF(y)=yx^{\prime}\}.$
@@
The first statement is proved in [DI24]. We show the second one. Let $(x,x^{\prime},y)\in{}^{\flat}\Sigma_{w}^{-m}$. By definition,
$y\in\mathbb{G}^{1}\mathbb{B}{\dot{w}}(c\sigma)^{2N-m}\mathbb{B}(c\sigma)^{-m+2N}=\mathbb{U}\mathbb{T}\mathbb{U}^{-,1}{\dot{w}}.$
@@
It follows that $z_{1}={}^{F(\tau)}h_{+}$, $L(\tau)=h_{0}^{-1}$ and $z_{3}=h_{-}$. Therefor,
$\Sigma_{w}^{i}=\{(\tau,z_{2},z_{4})\in\mathbb{T}\times(F\mathbb{U}\cap\mathbb{U}^{-})\times(F\mathbb{U}^{-,1}\cap\mathbb{U});L(\tau)={\rm pr}_{0}({}^{F(\tau)^{-1}}z_{2}z_{4})\},$
@@
$H_{*}(\Sigma_{w}^{2N-m})=H_{*}({}^{\flat}\Sigma_{w}^{2N-m})\cong H_{*}((\Sigma_{w}^{i})^{\mathbb{T}})=H_{*}({\dot{w}}\mathbb{T}^{F})$
@@
Let $P$ be the coweight be the coweight lattice of $\Phi$. If $\mu=0\in P/(1-c\sigma)P$, then $\Delta_{e}=\Delta$ and the statement is trivial. So we may assume that $X_{*}(T)/(1-c)X_{*}(T)\neq\{0\}$, which excludes the types ${}^{2}A_{n-1}$ ($n$ even), ${}^{3}D_{4}$, $E_{8}$, ${}^{2}E_{6}$, $F_{4}$, $G_{2}$. Then we will take a case-by-case analysis for the remaining types.
We adopt the labeling of Dynkin diagrams by positive integers as in [Hum72]. For $i\in\mathbb{Z}_{\geqslant 1}$. let $s_{i}$ and $\omega_{i}^{\vee}$ denote the corresponding simple reflection and fundamental coweight, respectively.
Case (1): $\Delta$ is of type $A_{n-1}$. Take $c=s_{1}s_{2}\cdots s_{n-1}$. Then we have $P/(1-c\sigma)P=\{0,\omega_{1}^{\vee},\omega_{2}^{\vee},\dots\omega_{n-1}^{\vee}\}$. Assume $\mu=\omega_{k}^{\vee}$ with $k\in\mathbb{Z}$. Let $m={\rm{gcd}}(k,n)\in\mathbb{Z}_{\geqslant 1}$. Then we take $I$ to be the complement of the sequence $I^{c}=(n/m,2n/m,\cdots,(m-1)n/m)$.
@@
Case (5): $\Delta$ is of type $D_{n}$ with $n\geqslant 4$. Take $c=s_{1}s_{2}\cdots s_{n}$. Then $P/(1-c\sigma)P=\{0,\omega_{1}^{\vee},\omega_{n-1}^{\vee},\omega_{n}^{\vee}\}$. If $\mu=\omega_{1}^{\vee}$, take $I=(n-1,n)$. It remains to handle the case $\mu=\omega_{n-1}^{\vee}$ by symmetry. If $n$ is even, take $I=(1,3,\dots,n-3,4)$ if $4\mid n$ and $I=(1,3,\dots,n-3,n-1)$ if $4\nmid n$. If $n$ is odd, take $I=(1,3,\dots,n-3,n-4,\dots,n-2,n)$.

### v2

@@ 3.2. An extension of action
We define an action of $\mathbb{T}^{F}\times\mathbb{T}^{F}$ on $\hat{\Sigma}_{w}^{i}$ by
@@
$H_{w,\chi}=\{(t,t^{\prime})\in\mathbb{T}\times\mathbb{T};w^{-1}t^{-1}F(t)w={t^{\prime}}^{-1}F(t^{\prime})\in\operatorname{Im}(\chi)\}.$
@@
$N_{F}^{F^{N_{0}}}:\mathbb{T}\mathbb{T},\quad ttF(t)\cdots F^{N_{0}-1}(t).$
@@ 3.3. Handling $\Sigma_{w}^{0}$ for $w\in W_{e}\setminus W_{e}^{c\sigma}$
$(\hat{Y}_{w,u}^{i})^{\mathbb{G}_{m}}\subseteq\{1\}\times\{1\}\times\{1\}\times\mathbb{T}\times\{1\}\times\{1\}.$
@@ 3.4. Handling $\Sigma_{w}^{0}$ for $w\in W_{e}^{c\sigma}$
${}^{\flat}\Sigma_{w}^{i}=\{(x,x^{\prime},y)\in(F\mathbb{U}\cap\mathbb{U}^{-})\times F^{i}(F\mathbb{U}\cap\mathbb{U}^{-})\times(\mathbb{B}{\dot{w}}\mathbb{G}^{1}F^{i}\mathbb{B});xF(y)=yx^{\prime}\}.$
@@
The first equality is proved in [DI24]. We show the last two equalities. Let $(x,x^{\prime},y)\in{}^{\flat}\Sigma_{w}^{2N-m}$. By definition,
$y\in\mathbb{G}^{1}\mathbb{B}{\dot{w}}(c\sigma)^{2N-m}\mathbb{B}(c\sigma)^{m-2N}=\mathbb{U}\mathbb{T}\mathbb{U}^{-,1}{\dot{w}}.$
@@
It follows that $z_{1}={}^{F(\tau)}h_{+}^{-1}$, $L(\tau)=h_{0}^{-1}$ and $z_{3}=h_{-}$. Therefore,
${}^{\flat}\Sigma_{w}^{2N-m}=\{(\tau,z_{2},z_{4})\in\mathbb{T}\times(F\mathbb{U}\cap\mathbb{U}^{-})\times(F\mathbb{U}^{-,1}\cap\mathbb{U});L(\tau)={\rm pr}_{0}({}^{F(\tau)^{-1}}z_{2}z_{4}^{-1})\},$
@@
$H_{*}(\Sigma_{w}^{2N-m})=H_{*}({}^{\flat}\Sigma_{w}^{2N-m})=H_{*}(({}^{\flat}\Sigma_{w}^{2N-m})^{\mathbb{T}})=H_{*}({\dot{w}}\mathbb{T}^{F})$
@@
Let $P$ be the coweight lattice of $\Phi$. If $\mu=0\in P/(1-c\sigma)P$, then $\Delta_{e}=\Delta$ and the statement is trivial. So we may assume that $P/(1-c)P\neq\{0\}$, which excludes the types ${}^{2}A_{n-1}$ ($n$ odd), ${}^{2}D_{n}$, ${}^{3}D_{4}$, $E_{8}$, ${}^{2}E_{6}$, $F_{4}$, $G_{2}$. Then we will take a case-by-case analysis for the remaining types.
We adopt the labelling of Dynkin diagrams by positive integers as in [Hum72]. For $i\in\mathbb{Z}_{\geqslant 1}$. let $s_{i}$ and $\omega_{i}^{\vee}$ denote the corresponding simple reflection and fundamental coweight, respectively.
Case (1): $\Delta$ is of type $A_{n-1}$. Take $c=s_{1}s_{2}\cdots s_{n-1}$. Then we have $P/(1-c\sigma)P=\{0,\omega_{1}^{\vee},\omega_{2}^{\vee},\dots\omega_{n-1}^{\vee}\}$. Assume $\mu=\omega_{k}^{\vee}$ with $k\in\mathbb{Z}$. Let $m={\rm{gcd}}(k,n)\in\mathbb{Z}_{\geqslant 1}$. Then we take $I$ to be the complement of the sequence $I^{\prime}=(n/m,2n/m,\cdots,(m-1)n/m)$.
@@
Case (5): $\Delta$ is of type $D_{n}$ with $n\geqslant 4$. Take $c=s_{1}s_{2}\cdots s_{n}$. Then $P/(1-c\sigma)P=\{0,\omega_{1}^{\vee},\omega_{n-1}^{\vee},\omega_{n}^{\vee}\}$. If $\mu=\omega_{1}^{\vee}$, take $I=(n-1,n)$. It remains to handle the case $\mu=\omega_{n-1}^{\vee}$ by symmetry. If $n$ is even, take $I=(1,3,\dots,n-3,4)$ if $4\mid n$ and $I=(1,3,\dots,n-3,n-1)$ if $4\nmid n$. If $n$ is odd, take $I=(1,3,\dots,n-4,\dots,n-2,n)$.